Water can be decomposed into H2 and O2 via electrolysis. How many moles of H2 could theoretically be obtained from 87.5 g of water?
1
Expert's answer
2017-10-13T15:11:07-0400
H 2 O = H 2 + ½ O 2
n(H 2 ) = n(H 2 O) = m(H 2 O)/M(H 2 O) = 87.5/18 = 4.86 (mol)
Answer: 4.86mol of H 2 could theoretically be obtained from 87.5 g of water.